• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name H. Ibrahim

    Log entry time 22:18:17 on November14,2002

    Entry number 90105

    keyword=Target Problems

    At about 20:30 I noticed that the temp of loop 2 (LD2) is decreasing constantly and the target alarm handler is beeping continously with red colors everywhere. I asked the MCC to take the beam until we investigate the problem. I used the aux heater when the hardware target alarm went off because of low temp and then rebooted the IOC. During the time of rebooting, the temp of loop 2 was changing rapidly up and down without any clear reason even after the target controls came back alive. At that point I paged the target on call (J.P. Chen) to help us. J. P. called us back and told me to turn off the aux heater and to check up everything on the target controls, but that did not help it and then he told me to turn off the beam compensation mode and to change the PID mode from PID to MAN and then start changing slowly the high power heater manually to set the temp at 22 K. After that I had to turn on again the beam compensation and change the PID mode to PID when we got 22.00 K manually. After doing all this the loop 2 temp bacame stable and we were ready to ask the MCC to give us back the beam at 21:23. Beam came back at 21:35 and now everything looks good.


    FIGURE 1

    FIGURE 2